Displaying report 1-1 of 1.
Reports until 17:02, Tuesday 15 October 2024
Equipment General
Torrey Cullen - posted 17:02, Tuesday 15 October 2024 - last comment - 11:19, Wednesday 16 October 2024(11907)
Optics move to B111

[Ian, Daniel, Jeff, Torrey]

Moving of the optics sleds has been completed without any hiccups. Particle count pre moving was 0/0/0 and 20 minutes after things have settled 0/0/0. The mylar is still over the cavities, we will unwrap when we plan on turning on lasers. Pictures attached.

Please note that the fans in the mobil cleans rooms in B111 should never be turned off now.

Images attached to this report
Comments related to this report
Chris Stoughton - 08:43, Wednesday 16 October 2024 (11908)

Looks great.  Can you remind me what the units are on the particle counter?  What are the three particle sizes, and is it calibrated in terms of particles per volume?

Daniel Grass - 11:19, Wednesday 16 October 2024 (11909)

x/x/x is the 0.3 um, 0.5 um, and 1.0 um cumulative count over 1 minute.

Displaying report 1-1 of 1.